# Can Large Language Models Capture Video Game Engagement?

### v1

@@ Abstract
Can out-of-the-box pretrained Large Language Models (LLMs) detect human affect successfully when observing a video? To address this question, for the first time, we evaluate comprehensively the capacity of popular LLMs to annotate and successfully predict continuous affect annotations of videos when prompted by a sequence of text and video frames in a multimodal fashion. Particularly in this paper, we test LLMs’ ability to correctly label changes of in-game engagement in 80 minutes of annotated videogame footage from 20 first-person shooter games of the GameVibe corpus. We run over 2,400 experiments to investigate the impact of LLM architecture, model size, input modality, prompting strategy, and ground truth processing method on engagement prediction. Our findings suggest that while LLMs rightfully claim human-like performance across multiple domains, they generally fall behind capturing continuous experience annotations provided by humans. We examine some of the underlying causes for the relatively poor overall performance, highlight the cases where LLMs exceed expectations, and draw a roadmap for the further exploration of automated emotion labelling via LLMs.
@@ I Introduction
The use of autoregressive modelling and large pretrained models such as Large Language Models (LLMs) is currently dominating AI research. LLMs have demonstrated unprecedented advances in language translation, code generation, problem solving, and AI-based assistance among many other downstream tasks [1]. Given their versatility and efficiency compared to earlier autoregressive models, one might even argue that the current capabilities of LLMs are endless as long as a problem and its corresponding solution(s) are represented as text. Meanwhile, the recent applications of LLMs within affective computing largely consider text-based affect modelling tasks such as LLM-based sentiment analysis [2, 3, 4]. The automatic labelling of affect based on time-continuous visual input remains largely unexplored [4], however, as the handful of studies available rely on still images [5, 6].
Motivated by the aforementioned lack of studies this paper introduces the first comprehensive evaluation of LLMs tasked to predict time-continuous affect labels from videos. In this initial evaluation we let LLMs observe gameplay videos as we prompt them with textual information of what they observe, and ask them to label the viewer engagement on those videos. We chose games as the domain of our study since they can act as rich elicitors of emotions and can offer a wide range of dynamic scenes and stimuli, varying from intense player actions to less intense game-world exploration. Even though LLMs have been used in a series of diverse tasks within the domain of videogames—both in academic studies [7, 8] and industrial applications such as AI Dungeon (Latitude, 2019), AI People (GoodAI, 2025) and Infinite Craft [^1] https://neal.fun/infinite-craft/—the capacity of these foundation models as predictors of player experience has not been investigated yet.
@@
Figure 2: Overview of the evaluation experiments presented in this study. Independently of experimental setting, the downstream task is engagement prediction formulated as a binary preference. We use a combination of text prompts and/or video frames as input and task the LLMs to label engagement. To evaluate the models, we compare the generated labels to the ground truth labels from the annotated GameVibe corpus (see Section III). All LLMs are prompted with a Chain-of-Thought (CoT) strategy. In the Text Input setup, the input for the downstream task is text descriptions (see Section IV-C) whereas in the Multimodal settings, the input contains both images and text prompts (see Section IV-B). In the few-shot experiments we generate reasoning examples based on ground truth evaluations. The examples are given to the LLM in addition to the base CoT prompt and the images (see Section IV-D). In all experimental settings we generate a description, comparison, reasoning, and a decision relating to an increase or decrease in engagement. We parse these outputs to derive the final binary engagement evaluation. Descriptions are only generated in the Multimodal Input settings.
@@
We employ LLMs as autonomous player experience annotators and present a thorough evaluation of their capacity to predict player experience in one-shot and few-shot fashions. Specifically, we compare state of the art foundation models from the LLaVA and GPT families against human annotated data of player engagement of the GameVibe dataset [9] (see Fig. 1). The dataset contains continuous engagement labels of gameplay videos across a variety of first-person shooter (FPS) games. We present selected results out of $2,440$ experimental settings in which we vary and test LLM model types, model sizes, prompting strategies, input types, and ground truth processing methods. Figure 2 shows a high level overview of our experimental setup followed in this study. We focus on videogame footage—one might come across on game streaming services such as Twitch [^2] https://www.twitch.tv—as input and viewer engagement as output.
@@
The novelty of this paper is two-fold. First, we investigate the capacity of LLMs to accurately label affect in a time-continuous manner using videos as affect elicitors. Second, we present the first large set of evaluation experiments that lays the groundworks for LLM-based player experience prediction. Our experiments show the feasibility of leveraging LLMs for engagement prediction particularly on popular games with a rich online presence (such as ApexLegends, 2019). Our key findings suggest that a) text-based summarisation of frames and direct multimodal prompting do not impact LLM performance; b) LLM performance is largely dependent on the elicitor (i.e. different games in this study); c) the multimodal few-shot prompting strategy is the one that improves LLM performance the most; and d) scale matters. Specifically, the best results obtained are when we employ the GPT-4o model and we feed it with a few positive and negative multimodal examples of increasing or decreasing engagement (few-shot prompting). While this approach yields an average accuracy of $6\%$ over the baseline across games, the GPT-4o model is able to improve the baseline performance by up to $47\%$ in certain games.
@@ II-A LLMs for Affect Modelling
Amin et al. [19] have also conducted a study on the capabilities of GPT [20] on affective computing tasks. They have put forth a comprehensive series of experiments which included a similar pairwise preference classification task for engagement prediction to what we use in this paper. They showed that when it comes to subjective tasks with a high potential for disagreement between annotators, out-of-box LLMs, such as GPT struggle compared to architectures leveraging specialized supervised networks. In those experiments—focusing on a simple one-shot prompting strategy on text input—GPT barely surpassed the baseline. In contrast [19], we investigate multimodal, chain-of-thought, and few-shot strategies in visual-based engagement prediction tasks across multiple games, analysing where LLMs either struggle or flourish compared to baseline approaches.
@@ II-B LLMs in Games
The recent developments in LLM methods and technology brought unprecedented wide adoption of AI across multiple domains including law [21], healthcare [22], and education [23]. Advancements in transformer architectures [24], coupled with a rapid increase in dataset and parameter sizes [25] led to a new wave of algorithms with previously unseen capabilities to generate high-quality text. Starting with Bidirectional Encoder Representations from Transformers (BERT) [18] but eventually popularized with the release of Generative Pre-trained Transformers (GPT) [1, 26, 20], LLMs have largely been characterized as transformer-based models, using large amounts of parameters (in the 100 millions and billions), built on large amounts of data, generating text in an autoregressive manner—that is predicting future tokens based on prior data. More recently, LLMs have been expanded to handle new modalities beyond text, such as audio and images [13], making them a candidate for applications using multimodal content such as gameplay videos.
In the context of games, LLMs have been used to create game-playing agents [27, 28], commentators [29] game analytics [30, 31], AI directors and game masters [32, 33], content generators [34], and design assistants [35]. Beyond the academic setting, we are seeing considerable interest from industrial players as well, such as NVIDIA’s recent ACE small language models  [^4] https://developer.nvidia.com/ace for autonomously generating the behaviour and animation of NPCs. Gallotta et al. [7] offer a recent and thorough overview on how LLMs can be utilised in games. In their roadmap, they identify player modelling as one of the most promising, yet unexplored avenues for future research into LLMs and games. Whilst affect modelling research has demonstrated that LLMs can be effective predictors in tasks such as sentiment analysis [10], they are yet to be widely evaluated to modelling player experience in the context of games.
@@ II-C Player Affect Modelling
Player modelling is an active field within AI and games research [8] with a particular focus on methods that capture emotional and behavioural aspects of gameplay such as engagement [36], toxicity [37] and motivation [38]. Traditionally, the field has focused heavily on data aggregation [39] and pattern discovery [40, 38] of playing behaviours, but there has been a recent shift towards moment-to-moment predictive models of players [41, 36, 42, 43, 9]. The prevalent strategy of such modelling methods relies on the availability of continuous annotation traces, which are generally processed as interval data [44]. This allows for the treatment of the labelled data as absolute ratings such as player engagement levels or classes such as low and high game intensity [43, 41].
@@ III The GameVibe Corpus
This section gives a general overview of the GameVibe corpus used throughout all experiments presented in this paper followed by an outline of the preprocessing approach we adopted for the engagement labels in this study. While the dataset is introduced thoroughly in [9] in this section we highlight the main aspects of the dataset that are relevant to our experiments here.
@@ III-A Corpus Overview
GameVibe is organized into 4 sessions of 30 unique video clips of 1 minute each, with each video in a session annotated by the same set of 5 human annotators. The video clips were selected to contain a maximum of 15 seconds of non-gameplay content such as pause menus and cut scenes, and were sampled at 30 hertz with a resolution of $1280\times 720$ for modern titles and $541\times 650$ for older titles. Annotations were collected using the PAGAN annotation platform [52] and the RankTrace annotation tool [53] (see Fig. 3), with the videos presented to participants in random order to minimize habituation and ordering effects. In RankTrace, participants are exposed to stimuli and annotate in real-time by scrolling up or down on a mouse wheel in an unbounded manner to indicate increases and decreases of their labelled state, in this case viewer engagement. Participants of GameVibe were given the following definition of engagement prior to starting their annotation task:
@@
After a qualitative analysis of the dataset, we select 20 games from the GameVibe corpus to form GameVibe-LLM (see Table I). We discard 10 games that feature third-person segments, mix footage of menus and gameplay, have large mobile UI overlay, or include poor footage. We select one out of four sessions randomly for generating few-shot examples in the final experiments and we test performance on the remaining 3 sessions. To be able to fairly compare the performance of different setups, we exclude the selected session from the remaining of the experiments.
@@ III-B Engagement Data Pre-Processing
Our data preprocessing method closely follows common practices in affective computing and methods introduced in previous studies with GameVibe [54]. Thus, each annotation trace was resampled into three-second non-overlapping time windows using simple averaging. The videos were sampled at a similar rate to align the stimuli to the engagement traces provided by the participants. These traces were then processed into discrete ordinal signals by comparing pairs of consecutive time windows to determine whether engagement increased (1), decreased (-1) or remained stable (0) between the two time windows.
Based on preliminary experiments and findings from earlier studies [55, 54], we solely focus on the data points with changes in engagement. We thus removed any data where engagement remained stable and fix the time-window between frames at $3$ seconds. We select frames from the first minute of gameplay and extract $20$ videos per session. We discard the first comparison in each session (frame 0 to frame 1) because the very first frame of the videos lack necessary context for the viewer to provide meaningful a rating. This means we have $18$ comparisons per video. We select $3$ sessions from each game and after removing uncertain evaluations, we end up with around $2,000$ comparisons (around $33$ per video).
@@ IV Methodology
In this section we detail our chosen algorithms and the different prompting strategies we employ throughout our experiments. In the presented studies we evaluate the capacity of LLMs to correctly evaluate changes of engagement in gameplay videos. In particular we picked LLaVA and GPT-4o as our base LLMs under investigation (see Section IV-A). In all reported experiments the downstream task of the employed LLM is to label a change in engagement (increase or decrease) given two consecutive frames of a video. We evaluate the algorithm’s performance against the human labelled engagement data of GameVibe that we treat as our ground truth.
@@
To explore how different experimental setups affect LLM engagement predictability, we ran experiments both with Multimodal and Text Input. Figure 2 illustrates the overall strategy and the different experimental setting employed. In the Multimodal Input setting, the input for the algorithm is one or two images accompanied by a text-prompt describing the task. We detail the format of the multimodal input in Section IV-B. In the Text Input setting, instead, we provide text-based descriptions of two video frames as part of the text prompt. We describe the format of the text input in Section IV-C. Finally, we also study few-shot prompting, using multimodal input and we detail this process in Section IV-D along with our general prompting strategy.
IV-A Employed LLMs
As mentioned earlier, we employ the Large Language and Vision Assistant (LLaVA) [56] and the Generative Pre-trained Transformer (GPT) models for all reported experiments. This section outlines the reasons we select these two LMMs and details the specific algorithmic properties we used for each model.
IV-A1 LLaVA
LLaVA [57, 56] is an ensemble model connecting a vision encoder with an LLM. LLaVA uses Contrastive Language–Image Pre-training (CLIP) [58] as a vision encoder and Vicuna [59] as a language decoder. To train LLaVA, Liu et al. leveraged GPT4 to generate data on instruction following examples and trained their framework end-to-end to fuse vision and language input. The result is a robust model which is able to output text-descriptions and solve reasoning tasks based on image and text prompts combined. We have selected LLaVA because a) it is an open-source model with multimodal capabilities; and b) it is easily deployed in local environments. We run experiments with the $7$ billion (7b), $13$ billion (13b), and $34$ billion (34b) parameter version of the algorithm using the Ollama API [^5] https://ollama.com/.
IV-A2 GPT-4o

### v2

@@ Abstract
Can out-of-the-box pretrained Large Language Models (LLMs) detect human affect successfully when observing a video? To address this question, for the first time, we evaluate comprehensively the capacity of popular LLMs for successfully predicting continuous affect annotations of videos when prompted by a sequence of text and video frames in a multimodal fashion. In this paper, we test LLMs’ ability to correctly label changes of in-game engagement in 80 minutes of annotated videogame footage from 20 first-person shooter games of the GameVibe corpus. We run over 4,800 experiments to investigate the impact of LLM architecture, model size, input modality, prompting strategy, and ground truth processing method on engagement prediction. Our findings suggest that while LLMs rightfully claim human-like performance across multiple domains and able to outperform traditional machine learning baselines, they generally fall behind continuous experience annotations provided by humans. We examine some of the underlying causes for a fluctuating performance across games, highlight the cases where LLMs exceed expectations, and draw a roadmap for the further exploration of automated emotion labelling via LLMs.
@@ I Introduction
The use of autoregressive modelling and large pretrained models such as Large Language Models (LLMs) is currently dominating AI research. LLMs have demonstrated unprecedented advances in language translation, code generation, problem solving, and AI-based assistance among many other downstream tasks [59]. Given their versatility and efficiency compared to earlier autoregressive models, one might even argue that the current capabilities of LLMs are endless as long as a problem and its corresponding solution(s) are represented as text. Meanwhile, the recent applications of LLMs within affective computing largely consider text-based affect modelling tasks such as LLM-based sentiment analysis [87, 13, 88]. The automatic labelling of affect based on time-continuous visual input remains largely unexplored [88]. The handful of studies focusing on multimodal input, still largely rely on images [35, 80] or—in rare cases where we see video and/or audio input [92]—they focus on larger units, evaluating whole scenes instead of time-continuous labelling.
@@
Motivated by the aforementioned lack of studies, this paper introduces the first comprehensive evaluation of LLMs tasked to predict time-continuous affect labels from videos. We observe the ability of various LLMs to rank consecutive video frames in terms of viewer engagement. While engagement is generally not considered a well-defined emotional reaction, emotional engagement encompasses a broad range of measurable affective reactions—including evaluative judgments, facial changes, autonomic responses (such as skin conductance and heart rate) [18]—that strongly correlate with user preferences [7], and can be conceptualized as a high arousal, positive valence affective response under the Circumplex Model of affect [63]. Throughout this paper we use engagement as an affective measure that captures the complexity of a positive emotional experience [7] of multimedia viewing. We chose games as the domain of our study since they can act as rich elicitors of emotions and can offer a wide range of dynamic scenes and stimuli, varying from intense player actions to less intense game-world exploration. Even though LLMs have been used in a series of diverse tasks within the domain of videogames—both in academic studies [26, 84] and industrial applications such as AI Dungeon (Latitude, 2019), AI People (GoodAI, 2025) and Infinite Craft [^1] https://neal.fun/infinite-craft/—the capacity of these foundation models as predictors of player experience has not been investigated yet.
@@
Figure 2: Overview of the evaluation experiments presented in this study. Independently of experimental setting, the downstream task is engagement prediction formulated as a binary preference. We use a combination of text prompts and/or video frames as input and task the LLMs to label engagement. To evaluate the models, we compare the generated labels to the ground truth labels from the annotated GameVibe corpus (see Section III). All LLMs are prompted with a Chain-of-Thought (CoT) strategy. We experiment with three different setups for possible inputs: Multimodal One-Shot—based on image input; Text Descriptions—based on LLM-generated descriptions of single frames; and Multimodal Few-Shot—based on LLM-generated positive and negative examples of engagement evaluation based on the ground truth. In all experimental settings we generate a description, comparison, reasoning, and a decision relating to an increase or decrease in engagement. We parse these outputs to derive the final binary engagement evaluation. Descriptions are only generated in the Multimodal settings.
@@
We employ LLMs as autonomous player experience annotators and present a thorough evaluation of their capacity to predict player experience in one-shot and few-shot fashions. Specifically, we compare state of the art foundation models from the LLaVA and GPT families against human annotated data of viewer engagement of the GameVibe dataset [8] (see Fig. 1). The dataset contains continuous engagement labels of gameplay videos across a variety of first-person shooter (FPS) games. We present selected results out of $4,840$ experimental settings in which we vary and test LLM model types, model sizes, prompting strategies, input types, and ground truth processing methods across $20$ games. Figure 2 shows a high level overview of our experimental setup followed in this study. We focus on videogame footage—one might come across on game streaming services such as Twitch [^2] https://www.twitch.tv—as input and viewer engagement as output.
@@
The novelty of this paper is two-fold. First, we investigate the capacity of LLMs to accurately label affect in a time-continuous manner using videos as affect elicitors. Second, we present the first large set of evaluation experiments that lays the groundwork for LLM-based player experience prediction. Our experiments show the feasibility of leveraging LLMs for engagement prediction, particularly on popular games with a rich online presence (such as ApexLegends, 2019). Our key findings suggest that a) text-based summarisation of frames and direct multimodal prompting do not impact LLM performance; b) LLM performance is largely dependent on the elicitor (i.e. different games in this study); c) the multimodal few-shot prompting strategy is the one that improves LLM performance the most; and d) scale matters. Specifically, the best results obtained are when we employ the GPT-4o model and we feed it with a few positive and negative multimodal examples of increasing or decreasing engagement (few-shot prompting). While this approach yields an average accuracy of $6\%$ over a Zero Rule Classifier across games, the GPT-4o model is able to improve the baseline performance by up to $47\%$ in certain games. Even though this performance is far from consistent, all tested LLMs exceeded traditional machine learning baselines in every experiment, underscoring their promise for complex affect modelling tasks.
@@ II-A LLMs for Affect Modelling
Amin et al. [3] have also conducted a study on the capabilities of GPT [51] on affective computing tasks. They have put forth a comprehensive series of experiments which included a similar pairwise preference classification task for engagement prediction to what we use in this paper. They showed that when it comes to subjective tasks with a high potential for disagreement between annotators, out-of-box LLMs, such as GPT struggle compared to architectures leveraging specialized supervised networks. In those experiments—focusing on a simple one-shot prompting strategy on text input—GPT barely surpassed the baseline. In contrast to [3], we investigate multimodal, chain-of-thought, and few-shot strategies in visual-based engagement prediction tasks across multiple games, analysing where LLMs either struggle or flourish compared to baseline approaches.
While our work examines the capabilities of general-purpose LLMs without any task-specific fine-tuning, recent research has increasingly turned to building specialized multimodal architectures and datasets. In particular, HumanOmni [92] introduces a large-scale corpus of human-centric data and an adaptive multimodal fusion architecture fine-tuned for emotion recognition, facial expression description, and action understanding. Similarly, FaVChat [91] focuses on human-centric content, but narrows its scope to fine-grained facial video analysis. While such domain-specific approaches demonstrate clear performance gains, they rely on extensive pre-training over curated datasets. In contrast, here we evaluate general-purpose, out-of-the-box models without additional fine-tuning.
@@ II-B LLMs in Games
The recent developments in LLM methods and technology brought unprecedented wide adoption of AI across multiple domains including law [32], healthcare [50], and education [48]. Advancements in transformer architectures [73], coupled with a rapid increase in dataset and parameter sizes [31], led to a new wave of algorithms with previously unseen capabilities to generate high-quality text. Starting with Bidirectional Encoder Representations from Transformers (BERT) [22] and eventually popularized with the release of Generative Pre-trained Transformers (GPT) [59, 24, 51], LLMs have largely been characterized as transformer-based models, using large amounts of parameters (in the 100 millions and billions), built on large amounts of data, generating text in an autoregressive manner—that is predicting future tokens based on prior data. More recently, LLMs have been expanded to handle new modalities beyond text, such as audio and images [70], making them a candidate for applications using multimodal content such as gameplay videos.
In the context of games, LLMs have been used to create game-playing agents [72, 29], commentators [61] game analytics [75, 62], AI directors and game masters [85, 93], content generators [67], and design assistants [25]. Beyond the academic setting, we are seeing considerable interest from industrial players as well, such as NVIDIA’s recent ACE small language models  [^4] https://developer.nvidia.com/ace for autonomously generating the behaviour and animation of NPCs. Gallotta et al. [26] offer a recent and thorough overview on how LLMs can be utilised in games. In their roadmap, they identify player modelling as one of the most promising, yet unexplored avenues for future research into LLMs and games. Whilst affect modelling research has demonstrated that LLMs can be effective predictors in tasks such as sentiment analysis [41], they are yet to be widely evaluated for modelling player experience in the context of games.
@@ II-C Player Affect Modelling
Player modelling is an active field within AI and games research [84] with a particular focus on methods that capture emotional and behavioural aspects of gameplay. Studies into player affect modelling often focus on arousal and valence [63]—aiming to deconstruct this emotional experience [45]. However, under the same theoretical paradigm, we can also conceptualize more complex emotional patterns as a function of affective dimensions. Given the complexity of both game playing and game spectating [7, 19], multiple studies have focused on higher level constructs, such as fun [9] and engagement [43, 8, 57]. Engagement in both gameplay and spectatorship can be characterized as a high arousal state [30]—with a positive valence in contrast to anxiety or frustration [19].
Traditionally, the field has focused heavily on data aggregation [23] and pattern discovery [40, 42] of playing behaviours, toxicity [14] and motivation [42]; but there has been a recent shift towards moment-to-moment predictive models of players [39, 43, 45, 11, 8]. The prevalent strategy of such modelling methods relies on the availability of continuous annotation traces, which are generally processed as interval data [82]. This allows for the treatment of the labelled data as absolute ratings or classes such as low and high intensity [11, 39].
@@ III The GameVibe Corpus
This section gives a general overview of the GameVibe corpus used throughout all experiments presented in this paper followed by an outline of the preprocessing approach we adopted for the engagement labels in this study. While the dataset is introduced thoroughly in [8], in this section we highlight the main aspects of the dataset that are relevant to our experiments here.
@@ III-A Corpus Overview
Each video in the GameVibe dataset is $1$ minute long, with every video annotated by the same set of $5$ human annotators. All $30$ games in the dataset have $4$ play sessions. The video clips were selected to contain a maximum of $15$ seconds of non-gameplay content such as pause menus and cut scenes, and were sampled at $30$ hertz with a resolution of $1280\times 720$ for modern titles and $541\times 650$ for older titles. Annotations were collected using the PAGAN annotation platform [44] and the RankTrace annotation tool [37] (see Fig. 3), with the videos presented to participants in random order to minimize habituation and ordering effects. In RankTrace, participants are exposed to stimuli and annotate in real-time by scrolling up or down on a mouse wheel in an unbounded manner to indicate increases and decreases of their labelled state, in this case viewer engagement. Participants of GameVibe were given the following definition of engagement prior to starting their annotation task:
@@
After a qualitative analysis of the dataset, we select $20$ games from the GameVibe corpus to form GameVibe-LLM (see Table I). We discard 10 games that feature third-person segments, mix footage of menus and gameplay, have large mobile UI overlay, or include poor footage. We select one out of four play sessions for each game to generate Few-Shot examples in the final experiments. We test performance on the three remaining play sessions for every game. To be able to fairly compare the performance of different setups, we exclude sessions withheld for the Few-Shot examples from all experiments.
@@ III-B Engagement Data Pre-Processing
Our data preprocessing method closely follows common practices in affective computing and methods introduced in previous studies with GameVibe [56, 55]. Thus, each annotation trace was resampled into $3$-second non-overlapping time windows using simple averaging. The videos were sampled at a similar rate to align the stimuli to the engagement traces provided by the participants. We consider the task a pairwise Preference Learning problem and aim to capture the dynamic change between frames. These traces were then processed into discrete ordinal signals by comparing pairs of consecutive time windows to determine whether engagement increased or decreased between the two time windows. We discard the first comparison in each play session (frame 0 to frame 1) because the very first frame of the videos might lack necessary context. This yields $18$ comparisons per play session—with $4$ play sessions per game, across $20$ games—totalling $1,440$ comparisons.
@@ IV Methodology
In this section we detail our chosen algorithms and the different prompting strategies we employ throughout our experiments. In the presented studies we evaluate the capacity of LLMs to correctly evaluate changes of engagement in gameplay videos. In particular we picked LLaVA and GPT-4o as our base LLMs under investigation (see Section IV-B). In all reported experiments the downstream task of the employed LLM is to label a change in engagement (increase or decrease) given two consecutive frames of a video. We evaluate the algorithm’s performance against the human labelled engagement data of GameVibe that we treat as our ground truth.
@@
To explore how different experimental setups affect LLM engagement predictability, we ran experiments both with Multimodal and Text Descriptions. Figure 2 illustrates the overall strategy and the different experimental settings employed. In the Multimodal Input setting, the input for the algorithm is one or two images accompanied by a text-prompt describing the task. We detail the format of the multimodal input in Section IV-C. In the Text Descriptions setting, instead, we provide text-based descriptions of two video frames as part of the text prompt. We describe the format of the text input in Section IV-D. Finally, we also study few-shot prompting, using multimodal input and we detail this process in Section IV-E along with our general prompting strategy.
IV-A Baseline Algorithms
As a point of comparison to our employed LLMs, we test two baseline algorithms on the GameVibe-LLM corpus. Both baselines were trained using 4-fold cross-validation—i.e. leaving one session out—and using early stopping based on prediction accuracy in the validation set. Both models are trained via cross entropy loss and the Adam optimizer.
For the first baseline, we extracted visual embeddings from frames in the corpus using the pre-trained DinoV2 “gigantic” [53] model, resulting in each image being compressed into a 1536 element vector. Prior to passing images through the model, each channel was normalized using the mean and standard deviations from the ImageNet dataset [21], which was the training data for the backbone models. We then trained a simple MLP classifier to map these embeddings to the engagement labels. The network consists of 2 layers of 1024 and 512 neurons, respectively, and was trained for a maximum of 300 epochs.
Second, we test a convolutional neural network (CNN) classifier using the raw frames as input. The RGB frames were normalized and resized to $(128\times 128)$. The CNN consists of four convolutional blocks, each with a kernel size of 3, followed by two fully connected layers of size $25,088$ and 512, respectively, with a dropout layer (rate of $0.5$) in between.
IV-B Employed LLMs
We employ a selection of different LLMs to provide a comprehensive look on the general capabilites of these models. We sample popular models with multi-modal (multi-image) capabilities from theHuggingface [^5] https://huggingface.co/ open source repository. We focus on models with most downloads and select the following LLMs: Phi 3.5-Vision; Gemma 3; InternVL2.5; LLaVA-Onevision; Qwen2.5-VL; and GPT-4o. The selected open-source LLMs fall between $4$ and $8$ billion parameters. While the models can be considered only ”medium“ in size, their popularity in the community attests their widespread use. We selected GPT-4o as a benchmark for large stat-of-the-art closed-source LLMs. The parameter size of GPT-4o has not been publicly disclosed.
IV-B1 Phi-3.5 Vision
Phi-3.5 Vision (Phi 3.5) is a compact multimodal model from Microsoft, extending the Phi-3 series multimodal capabilities [1]. The architecture consists of a CLIP-based ViT [58] and transformer decoder. To support different resolutions, the system divides images to multiple blocks and concatenates them. The text and image tokens are interleaved in the input space of the LLM backbone. The model is pre-trained to predict text output, then further tuned using both Supervised Fine-Tuning and Direct Preference Optimization [60]. Despite its smaller size ($4.15$ billion parameters ), Phi 3.5 is designed for high efficiency and instruction-following performance, trained on both curated synthetic and human-annotated data during. The model is optimized for reasoning and captioning tasks and provides competitive accuracy with low resource requirements. For our experiments, we use instruction-tuned variant. [^6] https://huggingface.co/microsoft/Phi-3.5-vision-instruct
IV-B2 Gemma 3
Gemma 3 is a member of the third-generation of Gemma models developed by Google DeepMind [68]. Gemma 3 uses a decoder-only transformer architecture and supports both text and vision modalities. The model’s vision encoder is derived from the SigLIP family [86], using $896\times 896$ pixel fixed-resolution tiles. This vision encoder is shared by all Gemma models regardless of the parameter size of the LLM backbone. We chose the $4.3$ billion parameter instruction-tuned variant for our experiments. [^7] https://huggingface.co/google/gemma-3-4b-it Despite the smaller parameter size, Gemma 3 4b was trained on a large budget of just over $4$ trillion tokens to be able to compete with medium sized LLMs. The final model is fine-tuned using a combination of knowledge distillation via an instruction-tuning teacher network and a wide array of Reinforcement-Learning objectives—including weight averaged reward models, human feedback, and ground-truth rewards.
IV-B3 InternVL2.5
InternVL2.5 belongs to a family of robust open-source models. It was developed by OpenGVLab [15] to match the performance of commercial LLMs—such as GPT-4o [51] and Claude-3.5-Sonnet [4]. The model is based on the InternViT Vision Transformer a small 2-layer MLP projector an an LLM backbone. InternVL2.5 is trained iteratively through Native Multimodal Pre-Training, Supervised Fine-Tuning, and Mixed Preference Optimization (MPO) [76]. InternVL2.5 integrates visual and linguistic signals using a fine-grained alignment strategy and is trained on a large multimodal corpus including real-world image-text pairs and synthetic instruction-following data. The 2.5 version introduces improvements in alignment accuracy and spatial reasoning. We chose the $8$ billion parameters (8b) version with the aforementioned MPO fine-tuning, which offers a balance between scalability and performance and has demonstrated strong results on visual question answering. [^8] https://huggingface.co/OpenGVLab/InternVL2_5-8B-MPO
IV-B4 LLaVA-OneVision
LLaVA-OneVision (LLaVA-OV) is a multimodal model based on the architecture of LLaVA [36]. It uses a SigLIP-based vision encoder [86], a 2-layer MLP projector, and Qwen2 as the LLM backbone. The model is trained in a 3-step process, which includes Language-Image Alignment, High-Quality Knowledge Learning, and finally Visual Instruction Tuning [34]. LLaVA-OV was trained on a wide array of single image, multi-image and video tasks and demonstrates strong multimodal understanding. By incorporating the Qwen2 language model, it benefits from a stronger language prior compared to earlier Vicuna-based [16] versions [36]. The model is instruction-tuned to perform general-purpose visual reasoning tasks. We selected the $8$ billion parameter model [^9] https://huggingface.co/llava-hf/llava-onevision-qwen2-7b-ov-hf based on Qwen2 7b for our experiments.
IV-B5 Qwen2.5-VL
Qwen2.5-VL is a multimodal variant of Qwen2.5, developed by Alibaba Group [5]. Qwen2.5-VL uses a custom Vision Transformer (ViT) combined with a 2-layer Multi Layer Perceptron (MLP), which projects the vision features into that aligns with the text embeddings of the LLM backbone. This method allows Qwen2.5-VL to process multiple images of varying resolution (or videos of varying length). The model is trained for structured multimodal reasoning, using spatial and temporal grounding via Multimodal Rotary Position Embedding—introduced in [74]. We use the 7b version of the instruction-tuned variant, which—despite the name—has $8.29$ billion (8.3b) parameters [^10] https://huggingface.co/Qwen/Qwen2.5-VL-7B-Instruct for our experiments.
IV-B6 GPT-4o